Find a number such that one-half of the number is 12 less than two-thirds of the number.
Answer by rfer(16322)

1/2x=2/3x-12
1/2x-2/3x=-12
3/6x-4/6x=-12
-1/6x=-12
x=-12/(-1/6)
x=-12*-6/1
x=72
--------------
1/2*72=2/3*72-12
36=36